# A Comprehensive Multi-Vocal Empirical Study of ML Cloud Service Misuses

### v1

@@ 4.2. Analysis of the Survey Results
Survey Participants: We reached a diverse group of 50 practitioners who actively participated in ML service-based projects, representing a wide range of roles within the field. Specifically, as shown in Figure 4, 34% of the participants identified as ML engineers, another 34% as data scientists, and 16% as software engineers. The remaining 16% of the participants held various other roles, including cloud engineers, data engineers, and software architects, highlighting the breadth of expertise and perspectives involved in the survey. In terms of experience, our participants also reflected considerable diversity. The majority reported having between one to five years of experience in AI (56%), software development (52%), and cloud computing (42%). Furthermore, 16% of the practitioners had less than one year of experience in both AI and software development, and 42% had less than one year of experience in cloud computing. This variety in experience levels further underscores the diversity of the participants, encompassing both emerging professionals and those with more established expertise in ML service-based systems.
Figure 4. Roles of Survey Participants
@@
Figure 5. Projects Distribution
Figure 6. Projects Scale
We asked practitioners about the frequency of using ML services. Figure 7 indicates that these services are commonly used in the industry, with 52% reporting monthly to daily usage, whereas only 16% use them once a year or less. This frequent use highlights the necessity for best practices. We also asked about which services are used most often and found Azure ML services, Amazon ML services, and Google ML services are at the top. Surprisingly, only 10% of practitioners use IBM Watson, and 6% use services from Oracle or others. Practitioners interact with ML services using various methods: 70% use APIs, 66% use Command Lines (CLI) and User Interfaces (UI), and 62% use SDKs, as shown in Figure 8. A preference for APIs is noted due to their integration ease, while the high usage of CLIs and UIs suggests a balance between technical and non-technical engagement.
We should note that diversity in (1) project sizes based on ML services, (2) the degree of ML services adoption, (3) interaction methods, and (4) different cloud providers enhances the generalizability of our findings, despite the number of participants in our survey.
Figure 7. Frequency of ML Service Utilization
Figure 8. Methods of ML Service Utilization
ML Services Impact and Challenges: Practitioners highlighted both positive and negative impacts of ML services on their project development process. On the positive side, 82% of practitioners reported that adopting ML services in their industrial projects helped them save development time. Also, 68% found that their projects were more scalable when using ML services, and 56% noted an increase in the efficiency of the software development process, as remarked in Figure9. These findings underscore the critical role of ML services in enhancing system scalability and efficiency while reducing the time required for development.
Practitioners have identified several challenges associated with ML services. Over half (54%) reported integration difficulties, indicating ML services can complicate the incorporation of new components into existing systems. Resource management issues were highlighted by 46% of practitioners, who experienced problems with memory and computational resource allocation. Security concerns were also prevalent, with 42% of participants expressing worries about data security and privacy. These challenges emphasize the need for effective practices to mitigate the negative impacts of ML services. In addition, 77% of practitioners faced challenges at various stages of their projects. Specific difficulties included data collection and processing (42%), training (38%), testing (35%), deployment (50%), serving (38%), and monitoring (40%). These findings illustrate the widespread nature of the challenges related to ML services and the need for comprehensive strategies to address them.
Figure 9. Impact of ML Services
Figure 10. ML Service-Related Challenges
Figure 11. Frequency of ML service challenges across different stages of the ML development pipeline
Takeaway 1: To fully unlock the benefits offered by ML services, it is important to address the challenges related to integration, resource management, and security. Tackling these issues will not only make projects succeed, but also boost significant benefits like efficiency, scalability, and improved capabilities in project development.
Catalog Agreement Analysis: We asked practitioners to what extent they agree or disagree that the practices in our catalog represent misuses of ML services. Figure 12 illustrates the level of agreement for each identified misuse. We found that all practices were considered misuses, with agreement levels ranging from 52% to 80%. The highest consensus (80%) was on misuses such as “Choosing the wrong deployment endpoint”, “Overwriting existing ML APIs without versioning”, “Ignoring fairness evaluation” and “Misinterpreting output”. On the other hand, the lowest agreement (52%) was for “Not using Batch API for data processing”. This lower score may reflect that some practitioners did not see this as a misuse, particularly in scenarios where small datasets make batch processing unnecessary. Furthermore, some practitioners were neutral, potentially due to varying contexts and use cases influencing their perceptions of these practices as misuses.
Takeaway 2: There is a strong consensus on the significance of our ML service misuse catalog, with the majority of practitioners acknowledging its crucial impact. This highlights a broad recognition of practical challenges within the industry, emphasizing the urgent need for strategic interventions and solutions.
Figure 12. ML Service Misuses Agreement Level
Correlation between Experience Levels and Agreement on ML Service Misuses: Our analysis of the correlation between participants’ experience levels in AI development and the number of ML service misuses they agreed upon revealed a moderate positive and statistically significant correlation (48%) between experience levels and agreement on ML service misuses. In addition, we studied the correlation between the participants’ experience level and their agreement level with each misuse. In 15 out of 20 misuses, we found positive and statistically significant correlations ranging from 30% to 47%. The positive correlation indicates that as developers gain more experience in AI development, they are more likely to agree on potential misuses in ML services, enabling them to identify misuse scenarios more effectively.
ML Service Misuses Frequency: We asked practitioners how frequently they encounter or tend to introduce each of the ML service misuses listed in our catalog. Specifically, for each misuse, practitioners were asked to indicate whether they have encountered the misuse in most ML projects (more than 75% of projects), in some ML projects (less than 75% of projects), or never encountered the misuse. To ensure reliability, for each misuse, we only retained responses from practitioners who only agreed on the misuse. As illustrated in Figure 13, our findings reveal that ML service misuses are prevalent in the industry, with only 10% to 32% of practitioners reporting that they had never encountered the misuses studied in their projects. We also found that “Improper handling of ML API limits” has the highest percentage of being encountered in most projects (31%). This could be due to the poor documentation available for setting proper ML API limits from cloud providers.
In addition, “Inefficient data transmission” has the highest percentage of being encountered in some projects (70%). This may reflect the common challenge of using caching mechanisms, for example, to process redundant data transfers. “Not using Batch API for data processing” is also a common problem in some projects (68%). This reflects the lack of awareness regarding the advantages of batch processing, resulting in the adoption of less optimal data processing methods. On the other end, “Excluding algorithms in automated ML” has the highest percentage of never being encountered (32%), followed by “Calling the wrong ML service API” (26%). The above findings suggest that these specific misuses are less common in practice.
Figure 13. ML Service Misuses Frequency
Takeaway 3: Misuses of ML services are prevalent across various stages of project development, with the most common issue being the improper handling of ML API limits. This oversight can lead to significant performance bottlenecks and service interruptions, thus addressing it becomes crucial to ensure smooth operations.
Correlation between Experience Levels and Frequency of ML Service Misuses: In nine of 20 misuses, we found positive and statistically significant correlations ranging from 29% to 36% between the level of experience of participants and the frequency of encountering each ML service misuse. This suggests that more experienced practitioners are better at identifying and recognizing these issues. This could indicate that more practiced professionals have a deeper understanding of ML service-based systems, are more familiar with common pitfalls, or are exposed to complex projects where such misuses are more prevalent.
Reasons of ML Service Misuses: We asked the practitioners about the possible reasons behind the presence of ML service misuses in practice. We included reasons in the survey based on our literature review (nahar2022collaboration, ; nahar2023meta, ; bogner2021characterizing, ) and provided an “Other” option to allow participants to add additional reasons freely. As shown in Figure 14, the most prominent issues identified were a lack of understanding of ML service capabilities (58%) and lack of documentation (50%). These findings highlight the urgent need for better educational resources and comprehensive documentation from ML cloud service providers. Also, half of the practitioners pointed to a lack of awareness or understanding of best practices for using ML cloud services and poor code review processes (44%) as significant reasons for introducing ML service misuses. This underscores the importance of disseminating best practices and establishing robust code review mechanisms. Other mentioned factors included pressure to meet tight deadlines (34%), and the rapid evolution of ML cloud services, which complicates staying updated with the latest changes (30%). As shown in Figure 14, no “Other” reasons were mentioned, which highlights the wide coverage of the predefined reasons in the survey. These insights suggest that addressing ML service misuses will require a multi-faceted approach, including improved documentation, and better education on ML capabilities, dissemination of best practices, and enhancing resource availability and expertise.
Takeaway 4: ML service misuse mainly stems from a lack of understanding of their capabilities, insufficient documentation, and limited awareness of best practices, which could lead to inefficiencies and poor performance. Hence, it is crucial to foster greater awareness and provide clear guidance to ML practitioners in this context to ensure more reliable and efficient deployments.
Figure 14. Reported Reasons of ML Service Misuses
Mitigation of ML Service Misuses: We asked the practitioners for mitigation solutions to avoid introducing ML service misuses.As shown in Figure 15, practitioners highly recommend to stay updated with best practices of ML services and cloud providers standards, as mentioned by 66% of the survey participants. This ensures that practitioners will be aware of the latest guidelines, tools, and techniques, reducing the likelihood of introducing potential ML service misuses. Furthermore, 64% of practitioners emphasized the importance of ensuring thorough validation of the quality of data used by ML components and conducting regular model and system performance reviews (58%). Other recommended practices include implementing robust alerting systems for model monitoring (48%), effective collaboration with domain experts and stakeholders (48%), and active engagement in code peer reviews (48%). These strategies collectively, reduce the risk of introducing ML service misuses, thus enhancing the quality and reliability of ML service-based systems.
Takeaway 5: To effectively mitigate ML service misuse, ML engineers should: (1) continually align with the latest best practices and standards from ML services and cloud providers, (2) rigorously validate ML data to ensure quality, and (3) systematically perform regular performance evaluations to detect issues early. These steps help maintain reliability and model accuracy.
Figure 15. ML Service Misuses Mitigation
Table 1. Summary of the Developed Catalog
ML Development Misuse Name Example Recommendation Frequency% Agree%
Pipeline Stage Most Some
@@
5. Discussion and Recommendations
Towards a better understanding of ML service misuses. We present the largest catalog of 20 ML service misuses (summarized in Table 1), which we identified through a multi-vocal study and a survey of 50 practitioners. We identify 290 occurrences of misuses based on the analysis of a curated set of 87 ML service-based systems. Our study does not only map out these misuses in the literature and ML service-based systems, but also quantifies practitioners’ agreement with our catalog and the frequency of encountering these misuses in the industry. We should note that the disparity between frequency (e.g., 68%) and agreement (e.g., 52%) for certain misuses, such as “Not using batch APIs for data loading”, highlights that while frequency indicates how often participants encounter the issue, agreement reflects a shared understanding of its problematic nature. This gap suggests that observing a misuse does not guarantee consensus on its consequences, likely due to variations in participants’ interpretations, project contexts, or professional backgrounds. Also, our survey results indicate that a significant percentage of participants have frequently encountered specific misuse in their professional experience, such as “Improper Handling of ML API limits” (31%). However, our analysis of open-source projects revealed only a few occurrences (30) of this misuse, highlighting a notable discrepancy.
There could be several reasons for this gap. Firstly, survey participants may work with production-level ML systems, where misuses are more common but often remain undocumented in public repositories. In addition, many ML misuses occur in private workflows, deployment pipelines, or closed-source production environments, making them less visible to open-source communities, such as GitHub, which was the basis for our analysis. Future research should explore this disconnect by examining industry case studies or interviewing practitioners to understand how such misuses are managed in real-world applications. Integrating insights from practitioners with open-source analysis would provide a comprehensive view of ML service misuses and their broader implications.
Moreover, the Spearman’s correlation between the frequency of encountering ML service misuses in most projects and their occurrence in GitHub projects is 0.46 (p-value = 0.04), indicating a moderate, positive, statistically significant correlation. This implies that as the frequency of ML service misuses increases, their occurrence across GitHub projects also tends to rise. The statistical significance (p ¡ 0.05) confirms that this relationship is unlikely to be due to random chance. This finding suggests that certain antipatterns are not only widely recognized and analyzed within software development but are also commonly observed in real-world projects. This prevalence can potentially affect the software quality and maintainability of software.
Towards more awareness of ML service misuses. The positive correlation between the experience level and the frequency of encountering those misuses, highlights the importance of knowledge transfer from experienced practitioners to less experienced ones, emphasizing the need to raise awareness about the good practices of ML services usage and improve the overall quality of ML service-based systems. This study has significant implications for various stakeholders in the ML ecosystem, including software/ML engineers, ML cloud providers, and open-source platforms like GitHub.
    For practitioners, our catalog of ML service misuses offers valuable insights into common pitfalls and misconfigurations that can reduce the reliability of ML service-powered systems. With this comprehensive catalog, practitioners can better understand potential risks and proactively design and integrate ML services that align with best practices. For example, software engineers, who often serve as the bridge between model development and production environments, can leverage this catalog to mitigate risks related to fairness, scalability, and performance.
    For cloud ML service providers, our study underscores the importance of developing tools and platforms that facilitate the detection and prevention of ML service misuse. We observed that some ML service misuses are trivial and could easily be spotted by service providers (e.g., “Overwriting existing ML APIs without versioning”). Hence, such tools would help ensure that users of ML services have access to transparent, scalable, and high-performance systems.
    Platforms such as GitHub and other open-source communities can play a critical role in promoting collaboration to identify, address, and mitigate these misuses, driving the creation of shared resources that increase the performance of ML service-based systems. For instance, GitHub could integrate a tool that detects ML service misuses when practitioners submit a pull request or commit their work, helping to flag potential misuses early in the development process.
Overall, our findings trigger an alarm towards investing in more robust tooling, policies, and guidelines to safeguard the integrity and societal impact of ML cloud services in particular and ML based systems in general.

### v2

@@ 5.1.4. Resulting Catalog of ML Service Misuses
Survey Participants: Our survey reached a diverse group of 50 practitioners who actively participated in ML service-based projects, representing a wide range of roles within the field. Specifically, as shown in Figure LABEL:fig:participants, 34% of the participants identified as ML engineers, another 34% as data scientists, and 16% as software engineers. The remaining 14% of the participants held various other roles, including data engineers (6%), software architects (6%), and cloud engineers (4%), highlighting the breadth of expertise and perspectives involved in the survey. In terms of experience, our participants also reflected considerable diversity. The majority reported having between one and five years of experience across various fields: AI (56%), software development (52%), and cloud computing (42%). Moreover, 42% had less than one year of experience in cloud computing, while 16% of the practitioners had less than one year of experience in both AI and software development. This variety in experience levels further underscores the diversity of the participants, encompassing both emerging professionals and those with more established expertise in ML service-based systems.
Size of Projects: We asked practitioners regarding the scale of ML service-based projects they had participated in. We found that more than half (52%) had worked on more than five such projects, while 48% had worked on one to five projects. Among these projects, as shown in Figure LABEL:fig:size, 44% were considered small ($\leq 5,000$ LOC), 46% medium-sized, and only 10% large ($\geq 30,000$ LOC). We also asked practitioners about the frequency with which they used ML services. Figure LABEL:fig:fre indicates that these services are commonly used in the industry, with 52% of respondents reporting usage Very Often (daily), Often (weekly), or Sometimes (monthly), while only 16% reported using them Rarely (once a year or less). This frequent use highlights the need for best practices. We further asked practitioners about which services are used most often, and found that Azure ML services, Amazon ML services, and Google ML services are among the top. Surprisingly, only a small proportion of practitioners indicated the use of IBM Watson (10%) and Oracle or others (6%). Practitioners indicated that they interact with ML services using various methods: 70% use APIs, 66% use Command Line Interface (CLI) and User Interfaces (UI), and 62% use SDKs, as shown in Figure LABEL:fig:meth. A preference for APIs is noted due to their integration ease, while the high usage of CLIs and UIs suggests a balance between technical and non-technical engagement.
We should note that the diversity in project sizes based on ML services, the degree of ML services adoption, interaction methods, and different cloud providers all enhance the generalizability of our findings, despite the survey’s number of participants.
ML Services Impact and Challenges: Practitioners highlighted both positive and negative impacts of ML services on their project development process. On the positive side, 82% of practitioners reported that adopting ML services in their industrial projects helped them save development time. In addition, 68% found that their projects were more scalable when using ML services, and 56% noted an increase in the efficiency of the software development process, as noted in Figure LABEL:fig:impacts. These findings underscore the critical role of ML services in improving system scalability and efficiency while reducing the time required for development. Practitioners also identified several challenges associated with ML services, which we report in Figure LABEL:fig:challenges. More than half (54%) reported integration difficulties, indicating that ML services can complicate the incorporation of new components into existing systems. The resource management issues were highlighted by 46% of the practitioners, who experienced problems with memory and computational resource allocation. Security concerns were also prevalent, with 42% of the participants expressing concerns about data security and privacy. These challenges emphasize the need for effective practices to mitigate the negative impacts of ML services. In addition, 77% of practitioners faced challenges at various stages of their projects. As shown in Figure LABEL:fig:stages, specific difficulties included data collection and processing (42%), training (38%), testing (35%), deployment (50%), serving (38%), and monitoring (40%). These findings show that the widespread challenges related to ML services can contribute to or result in their misuse, underscoring the need for comprehensive strategies to mitigate such risks.
Takeaway 1: It is important to address challenges in integration, resource management, and security to fully unlock the benefits of ML services. Doing so not only ensures project success but also enhances efficiency, scalability, and development capabilities.
Catalog Agreement Analysis: We asked practitioners about the extent to which they agree or disagree on whether the practices in our catalog represent misuses of ML services. Figure LABEL:fig:agreement illustrates the level of agreement for each identified misuse. We found that all practices were considered misuses by practitioners, with agreement levels ranging from 52% to 80%. The highest consensus (80%) was on misuses such as “Choosing the wrong deployment endpoint”, “Overwriting existing ML APIs without versioning”, “Ignoring fairness evaluation” and “Misinterpreting output”. In contrast, the lowest consensus (52%) was for “Not using Batch API for data processing”. This lower score may reflect that some practitioners did not deem this issue to be a misuse, particularly in scenarios where small datasets make batch processing unnecessary. Furthermore, some practitioners were neutral, potentially due to varying contexts and use cases influencing their perceptions of these practices as misuses.
Correlation Between Experience Levels and Agreement on ML Service Misuses: This analysis aimed at understanding whether awareness of ML service misuses correlates with practitioner expertise. This is particularly crucial because if misuse awareness is primarily lacking among less experienced practitioners, then targeted training and best-practice interventions should focus on this group. In contrast, if the correlation is weak or absent, then awareness-building efforts may need to be directed throughout the practitioner population. Our results on the correlation between participants’ experience levels in AI development and the number of ML service misuses they agreed on revealed a moderate positive and statistically significant correlation (48%) between experience levels and agreement on ML service misuses. In addition, we analyzed the correlation between the experience level of participants and their agreement level with each misuse. In 15 out of 20 misuses, we found positive and statistically significant correlations ranging from 30% to 47%. The positive correlation indicates that as developers gain more experience in AI development, they are more likely to agree on potential misuses in ML services, enabling them to recognize misuse scenarios more effectively. These findings emphasize the critical role of practitioner expertise in identifying misuse scenarios. While experienced practitioners are generally more capable of detecting and recognizing potential misuses, less experienced practitioners may benefit from targeted training and increased awareness of best and worst practices in ML service usage. However, the lack of strong correlation between expertise and certain misuses (e.g., “Inefficient data transmission”, “Avoiding parallel training experiments”, “Using suboptimal evaluation metrics”) suggests that some issues are less intuitive and thus require attention regardless of experience level.
Takeaway 2: There is strong consensus on the importance of our ML service misuse catalog, with most practitioners acknowledging its significant impact. More experienced professionals, in particular, show greater agreement, reflecting widespread awareness of industry challenges and the need for targeted solutions.
ML Service Misuses Frequency: We asked practitioners how frequently they encounter or introduce each of the ML service misuses listed in our catalog. Specifically, for each misuse, practitioners were asked to indicate whether they have encountered the misuse in most ML projects (i.e., more than 75% of projects), in some ML projects (i.e., less than 75% of projects), or never encountered the misuse. To ensure response reliability, we only retained answers where practitioners agreed on the presence of a misuse. As illustrated in Figure LABEL:fig:frequency, our findings reveal that ML service misuses are prevalent in the industry, with only 10% to 32% of practitioners reporting that they had never encountered such misuses in their projects. Notably, “Improper handling of ML API limits” was the most frequently observed misuse, with 31% of participants indicating its presence in most of their projects. This may be attributed to inadequate documentation from cloud providers on configuring appropriate API limits. In addition, “Inefficient data transmission” was the most frequently encountered misuse in some projects, reported by 70% of practitioners. This likely reflects common challenges such as the lack of caching mechanisms to avoid redundant data transfers. Similarly, “Not using Batch API for data processing” was reported in some projects by 68% of respondents, indicating limited awareness of the benefits of batch processing and a tendency toward less efficient data processing practices. On the other end, “Excluding algorithms in automated ML” had the highest percentage of never being encountered (32%), followed by “Calling the wrong ML service API” (26%). These findings suggest that such misuses are relatively less common in practice.
Takeaway 3: Misuses of ML services are prevalent throughout project development, with improper handling of ML API limits being the most common issue, often causing performance bottlenecks and service interruptions, which makes addressing it crucial for smooth operations.
Correlation Between Experience Levels and Frequency of ML Service Misuses: In nine of 20 misuses, we observed positive and statistically significant correlations (ranging from 29% to 36%) between participants’ experience levels and the frequency of encountering each ML service misuse. This suggests that more experienced practitioners are better at identifying and recognizing these issues. A possible explanation is that seasoned professionals have a deeper understanding of ML service-based systems, are more familiar with common pitfalls, or are exposed to more complex projects where such misuses are more likely to occur.
Reasons of ML Service Misuses: We asked practitioners about the possible reasons behind the presence of ML service misuses in practice. The survey options for possible reasons were derived from our literature review (nahar2022collaboration, ; nahar2023meta, ; bogner2021characterizing, ), and an Other option was included to allow participants to provide additional open-ended input. As shown in Figure LABEL:fig:reasons, the most prominent reasons were a lack of understanding of ML service capabilities (58%) and a lack of documentation (50%). These findings point to an urgent need for improved educational resources and more comprehensive documentation from ML cloud service providers. Moreover, half of the practitioners identified a lack of awareness or understanding of best practices for using ML cloud services, and 44% highlighted poor code review processes as key contributors to ML service misuse. This underscores the importance of disseminating best practices and establishing robust code review mechanisms. Other factors included pressure to meet tight deadlines (34%) and the rapid evolution of ML cloud services, which makes it challenging to stay up to date with the latest advancements (30%). Notably, no additional reasons were provided under the Other category, suggesting the predefined options offered broad coverage of the relevant reasons.
Furthermore, several of the identified factors were echoed in software engineering blogs and online developer discussions. In these sources, practitioners reported numerous practical challenges when working with ML cloud services (e.g., lack of understanding of ML service capabilities), which can often contribute to the introduction of ML service misuses. As a result, without sufficient guidance, developers may make incorrect assumptions or adopt suboptimal configurations. For instance, a Stack Overflow thread highlighted that users frequently encounter “TooManyRequests error” when interacting with Azure ML REST APIs (azureml_api_request_limits, ). This issue stems from unclear documentation regarding service request limits, forcing developers to implement their own rate-limiting strategies or rely on trial-and-error: “I could not find out how many concurrent requests are allowed by Azure Machine Learning batch endpoints, so I ended with a limit of 10 outgoing requests which solved the ”TooManyRequests” problem.” (azureml_api_request_limits, ). This example illustrates how the absence of clear documentation directly contributes to misuse or inefficient use of ML services. Similarly, a software quality engineer reported on Medium that “excessive time spent configuring build tools, dependency management, and deployment processes slows the development of ML service-based systems” (software_dev_problems, ). Such experiences help explain why practitioners often feel compelled to adopt ad-hoc solutions, rush deployments, or rely on trial-and-error approaches, mirroring our survey results where tight deadlines (34%) and rapid evolution of services (30%) were cited as reasons for misusing ML services. Also, Kastner et al. (kastner2025fairness, ) observed that “organizations often focus on short-term development goals and product success measures, whereas many fairness concerns relate to long-term outcomes, such as feedback loops, and avoiding rare disasters”. This supports the notion that economic and organizational pressures can drive technical decisions that compromise long-term quality, fairness, or responsible ML service usage.
These insights suggest that addressing ML service misuses requires a multi-faceted approach, including improved documentation, better education on ML capabilities, dissemination of best practices, and enhanced access to resources and expertise.
Takeaway 4: ML service misuse mainly stems from a lack of understanding of their capabilities, insufficient documentation, and limited awareness of best practices, which could lead to inefficiencies and poor performance, thus highlighting the need for better guidance and awareness among practitioners.
Mitigation of ML Service Misuses: We asked the practitioners for mitigation solutions to avoid introducing ML service misuses. As shown in Figure LABEL:fig:mitigation, the most commonly recommended approach was staying up to date with best practices for ML services and cloud provider standards, as mentioned by 66% of survey participants. This ensures that practitioners are aware of the latest guidelines, tools, and techniques, thereby reducing the likelihood of introducing potential misuses. In addition, 64% of practitioners emphasized the importance of thoroughly validating the quality of data used by ML components, while 58% highlighted the need for regular reviews of model and system performance. Other frequently recommended practices include implementing robust alerting systems for model monitoring (48%), collaborating effectively with domain experts and stakeholders (48%), and actively participating in peer code reviews (48%).
These strategies are also discussed in software engineering blogs and developer forums. For example, an ML engineer emphasized in a Medium blog the importance of “implementing a robust data monitoring system to track data quality and distribution over time, and setting up alerts or triggers to notify when significant drift is detected” (sohail2025datadrift, ). This aligns with the mitigation strategies identified in our survey, particularly conducting regular model performance reviews and implementing robust alerting systems for model monitoring. Furthermore, the same engineer emphasizes the importance of (1) “carefully preprocessing and cleaning data before training and deployment” and (2) “properly handling missing values, outliers, and inconsistencies to minimize the risk of data drift”. These practices reinforce our survey’s recommendation to ensure thorough validation of training data quality.
Moreover, a blog published by an AI-based knowledge management and automation company emphasized the importance of regular monitoring, stating: “(Step 3) Regular Monitoring–Continuously monitor fairness metrics as part of your model maintenance routine; Regular checks can help identify and address emerging biases over time” (zdrok2024fairness, ). This aligns with the mitigation strategy identified in our survey, which emphasizes the need for regular model performance reviews to proactively detect and address potential issues. Automated testing practices also support regular performance reviews. For example, a blog published by an AI-based testing solutions company highly recommends the use of Azure DevOps since it “facilitates lower manual error rates and faster and more reliable application delivery by integrating well-known test automation tools such as Selenium”. They also reported that it “allows for improved test coverage and early bug detection” (frugaltesting_azure_devops_testing, ). Broader reflections on documentation highlight its strategic importance: “A significant portion of a software project’s lifecycle (over 70%) is consumed by maintenance” which “underscores the essential role of documentation in facilitating effective software maintenance”, reported by an AI tech lead (nayeem_software_documentation, ). Developers also note that successful ML projects require strategic alignment across the organization, beyond technical execution. This supports our survey finding on the importance of close collaboration with domain experts and stakeholders. As noted in a Medium post, “building models is just one aspect of the process; it is crucial to involve feedback from users and cross-functional teams, incorporate new feature requests, and update initial hypotheses based on changing business needs.” (stahl_ai_ml_projects, ). Also, “to ensure the success of an AI project, it is crucial to have regular reviews and feedback sessions with stakeholders, including product teams and business leaders” (stahl_ai_ml_projects, ).
However, the practical implementation of these mitigation strategies in real-world ML workflows presents significant challenges. Staying current with best practices and cloud provider updates requires continuous effort, especially in rapidly evolving ecosystems. Likewise, validating data quality and conducting regular model reviews can be resource-intensive, demanding both organizational commitment and skilled personnel (lwakatare2021experiences, ; chen2022data, ; zhang2023automated, ). While automated testing and AI-assisted code review may alleviate some of this burden, their scope remains limited, and they require careful integration within existing CI/CD pipelines. To improve their effectiveness, there is a pressing need for dedicated tools that can automatically detect and refactor ML service misuses. Such tools would embed best practices at scale, providing actionable insights, enforce consistency, and reduce human error, thereby making mitigation strategies more practical and sustainable in production environments.
Takeaway 5: To effectively mitigate ML service misuse, ML engineers should: (1) continually align with the latest best practices and standards from ML services and cloud providers, (2) rigorously validate ML data to ensure quality, and (3) systematically perform regular performance evaluations to detect issues early.
5.3. RQ3. How comparable is the prevalence of ML service misuses in open-source projects to what is observed in industry practice?
To address this research question, we studied the prevalence of the 20 ML service misuses in GitHub projects compared to their frequency in industrial projects as reported by the survey practitioners (Figure LABEL:fig:frequency). Figure LABEL:fig:occ illustrates the prevalence of these misuses, along with the distribution of their occurrences across GitHub projects. As shown in Figure LABEL:fig:occ, “Ignoring fairness evaluation” is the most frequent misuse, appearing in over 60 projects, followed closely by “Ignoring monitoring for data drift” and “Inefficient data transmission”, which were found in approximately 50 and 40 projects, respectively. Other common misuses include “Not using batch APIs for data processing” and “Improper handling of ML API limits”, indicating that developers often struggle with data processing efficiency and API constraints. The figure also highlights less frequent but still notable misuses, such as “Misusing synchronous/asynchronous APIs”, Avoiding parallel training experiments”, and “Not using automatic hyperparameter tuning”, each found in fewer than 10 projects. Rare misuses, including “Disabling automatic rollbacks”, “Choosing the wrong deployment endpoint”, and “Excluding algorithms in AutoML”, occur in even fewer projects. Overall, the distribution suggests that, while some misuses are widespread and represent common problems in ML service development, others are rare but can still significantly impact model performance, reliability, and/or maintainability.
The Spearman’s correlation between the frequency of encountering ML service misuses in most industrial projects and their occurrence in GitHub projects is 0.46 (p-value = 0.04), indicating a moderate, positive, statistically significant correlation, confirming that this relationship is unlikely to be due to random chance. Specifically, this suggests that the distribution of ML service misuses identified in our manual analysis of GitHub projects closely mirrors those reported by practitioners. The prevalence of such misuses in real-world projects reinforces the need for increased awareness, improved best practices, and the development of automated tools to detect, prevent, and mitigate these issues. Addressing these misuses proactively could significantly enhance the reliability, efficiency, and cost-effectiveness of ML service-based systems in both enterprise and open-source domains. However, some discrepancies should be noted. Our survey results indicate that a significant percentage of participants have frequently encountered specific misuse in their professional experience, such as “Improper Handling of ML API limits” (31%). Yet, our analysis of open-source projects revealed that this misuse occurred only in 10% of the cases (30 out of 290), highlighting a notable gap. Several factors may explain this discrepancy. First, survey participants often work with production-level ML systems where misuses are more common but remain undocumented in public repositories. In addition, many ML misuses occur within private workflows, deployment pipelines, or closed-source production environments, making them less visible to open-source communities, such as GitHub, which formed the basis of our analysis. Also, the similar prevalence of ML service misuses in both open-source and industrial projects such as “Ignoring fairness evaluation” and “Ignoring monitoring for data drift” may partly stem from developers working with relatively stable or benchmark datasets, where changes in data distribution are perceived as unlikely. For example, when using state-of-the-art or well-established datasets, developers may feel confident enough to consciously deprioritize or omit monitoring mechanisms, which could explain why these misuses persist in practice. Prior work also reported that data drift monitoring is sometimes ignored when datasets are assumed to be neutral or lack protected attributes, leading developers to consider drift monitoring irrelevant in such contexts (polyzotis2019data, ).
Also, some practitioners reported having no systematic processes or tools to ensure fairness, instead relying on ad hoc detection by developers who “spot an issue that looks like a fairness issue to them, and then they talk with each other about it, and then find some specific solution to it” rather than employing structured and automated approaches (madaio2020co, ). Some teams even experimented with inferring demographics from indirect features to enable fairness checks, but expressed concerns that such indicators might themselves “introduce undesirable biases, introducing a need to audit the auditing tool” (holstein2019improving, ). Similarly, practitioners working on service or consulting tasks described challenges “relating to datasets outside of their control and about monitoring deployment contexts or fairness criteria after a system has been handed off to a customer” (madaio2020co, ). Hence, once a system is deployed, teams often lose visibility into the data and model usage, complicating fairness maintenance and causing issues to surface long after deployment. Together, these observations help explain why both “Ignoring fairness evaluation” and “Ignoring monitoring for data drift” remain prevalent: developers either deprioritize these tasks under assumptions of dataset stability or neutrality, or lack automated mechanisms and post-deployment visibility to consistently perform them. Future research should explore this disconnect by conducting industry case studies or practitioner interviews to understand how such misuses are managed in real-world applications. Integrating insights from practitioners with open-source analysis would provide a comprehensive understanding of ML service misuses and their broader implications.
Takeaway 6: Common ML service misuses, like ignoring fairness evaluation and data drift monitoring, appear similarly in both open-source projects and industry. However, some misuses, like improper handling of ML API limits, are more frequent in industry, reflecting differences between open-source and production environments.
6. Discussion and Recommendations
Towards a better understanding of ML service misuses. We present the largest catalog of 20 ML service misuses (summarized in Table LABEL:tab:catalog_summary), which we identified through a multi-vocal study and confirmed through a survey of 50 practitioners. We identified 290 occurrences of misuses based on the analysis of a curated set of 377 ML service-based systems. Our study not only maps out these misuses in the literature and ML service-based systems, but also quantifies practitioners’ agreement with our catalog and the frequency of encountering these misuses in the industry. Moreover, ML service misuses can create technical debt by embedding hidden complexities and maintenance challenges into the system. For example, ignoring data drift monitoring can cause ML service–based systems to make predictions on operational data that has shifted from the training distribution. Over time, this leads to performance degradation that is difficult to trace to its root cause, risking teams to implement repeated short-term fixes instead of addressing the underlying problem. These accumulated issues increase system complexity, elevate operational costs, and hinder the system’s ability to evolve or extend without costly refactoring.
@@ 6. Discussion and Recommendations
[…]
@@
While our survey focused on a subset of 20 validated misuses, we acknowledge that practitioner input on the remaining 14 misuses that we identified in our multivocal study could provide valuable additional insights. Exploring these misuses in future work represents a promising avenue for expanding our catalog and validating their relevance in real-world settings. Also, we should note that the disparity between frequency (e.g., 68%) and agreement (e.g., 52%) for certain misuses, such as “Not using batch APIs for data processing”, highlights that while frequency indicates how often participants encounter the issue, agreement reflects a shared understanding of its problematic nature. This gap suggests that observing a misuse does not guarantee consensus on its consequences, likely due to variations in participants’ interpretations, project contexts, or professional backgrounds. In addition, future research could take a reverse approach by first collecting misuses observed by practitioners in their projects and then examining their presence in open-source projects. This could reveal new, practice-driven misuse patterns and help refine specification approaches to better capture issues that arise in real-world systems. Moreover, a deeper analysis of our results segmented by cloud providers could offer valuable insights into potential differences in misuse patterns across platforms. Unfortunately, the survey data collected in this study did not include sufficient information to perform such an analysis. Future work should explore this aspect, as it could help uncover cloud-specific challenges and inform more targeted recommendations for practitioners.